# Demonstration Organization via Logit Separability: Advancing In-Context Learning through Multiple Class-Related Words

### v4

@@ 4.4.2 Effectiveness of Label Organization
In Table 1, incorporating multiple class-related words significantly boosts ICL performance across all models. We assess our label selection and ordering method by comparing it against a random organization of class-related words. In Table 5, the observed decreased accuracy in ‘+MLabels-Random’ underscores the superiority of our systematic approach.
@@
Table 5: Accuracy (%) comparison of our organization method with random organization in $P_{r}$ (same N as in Table 1).
@@ 4.5 Scalability to Larger LLMs
Table 6 presents the accuracy results of LICL implemented on LLaMA3-8b and LLaMA2-13b to further validate the effectiveness and applicability of our method on larger language models. The insertion of multiple class-related words led to an increase in accuracy. LICL achieved a 3.42% accuracy improvement on LLaMA3-8b and a 2.32% improvement on LLaMA2-13b. Specifically, for the IMDB dataset on LLaMA2-13b, despite significant sample information loss compared to other methods, our approach maintained strong performance. Details of the evaluation of initial class-related word updates can be found in Appx.G.2.
@@
4.6 Inserted Words’ General Applicability
To address potential concerns about overfitting when selecting multiple class-related words from a single dataset, we implemented cross-dataset evaluations. We evaluated the efficacy of using two class-related words (N in most binary tasks in our study) selected from the SST2 dataset across the CR and IMDB datasets, as detailed in Table 7.
@@ 4.6 Inserted Words’ General Applicability
+MLabels (from same dataset) 95.15 95.60 65.96 73.40
@@
In LLaMA2-7b, performance improvements on the CR may be attributed to similar sub-tokens found in both SST2 and CR (e.g., " unhealthy", " unfair", " good"). No performance decrease is observed in IMDB. In GPT2-xl, despite initial differences in the positive sentiment words across datasets, the class-related words from SST2 still enhanced performance in both CR and IMDB. These results suggest that the selected class-related words exhibit a level of generality, enhancing ICL performance across various datasets with similar labels.
@@ 4.7 Further Analyses
More in-depth analyses are in Appx.H, including pool filtering effectiveness and its ablation study (H.1), label bias evaluation (H.2), the impact of class-related words on predictions whether or not they appear in demonstrations (H.3), the effects of label-balanced demonstration in LICL (H.4), inference cost (H.5), and sample-label distribution visualizations (H.6).
@@ 5 Conclusion
In this paper, we introduce logit separability, a criterion for measuring the class-related clarity of samples and labels based on their logit value distributions, assessing whether a sample is easily predictable by its true label and whether a label performs well across multiple samples of the same class. To improve ICL with clearer class-related guidance, we propose LICL, a logit separability-based method for organizing demonstrations. Besides, to enrich ICL with broader label information, we present a novel label utilization method by forming sample-multiple-label pairs through the insertion of multiple class-related words, based on the logit distribution of selected samples. Experimental results confirm that this combined approach significantly enhances ICL classification performance, yielding superior outcomes.
@@ A.2 Label Effectiveness in ICL
A.2.1 Single Class-Related Word
This study establishes four distinct label sets for each dataset, utilizing identical samples to form the sample-label pairs in 1-shot ICL, as detailed in Table 8. The reported results represent the average accuracy obtained from five repeated experiments, conducted with seeds 42, 43, 44, 45, and 46 in sample selection during the 1-shot demonstrations. As shown in Fig.5, in LLaMA2-7b, the maximum 49.20% accuracy difference (TREC), and the highest 11.08% (SST2) standard deviation are observed. We also evaluate the label effectiveness in GPT2-xl under the same experimental conditions, with results shown in Fig.6. The results are similar to those in LLaMA2-7b, with a maximum accuracy difference of 48.20% (TREC) and a maximum standard deviation of 10.52% (SST2). These findings indicate that label selection contributes to both ICL accuracy and robustness.
@@ A.2.1 Single Class-Related Word
A.2.2 Multiple class-related words
@@ Appendix G Enhancement via Initial Class-Related Word Updates
Table 13 summarizes the effectiveness of initial class-related word updates in 5-shot experiments on the AMAN and TREC datasets, with word update information illustrated in Sec.4.2.
@@ H.4 Case Study: Effectiveness of Label Balance in LICL
The impact of label-unbalanced demonstrations varies between the two models. For LLaMA2-7b, a large-size language model, the negative effects of missing label information are less marked, with an average accuracy decrease of 3.37%. The most significant accuracy drops in TREC by 10.40%. The leverage of multiple class-related words mitigated the performance loss, reducing the average accuracy decline to 1.93% in ‘+MLabels_LW’ compared to a label-balanced setting. In contrast, GPT2-xl, a smaller model with 1.5 billion parameters, experienced a marked performance decline under unbalanced conditions, averaging a 12.89% decrease in accuracy. Specifically, SST2 and AGNews experienced over 34% declines. Interestingly, CR and ISEAR demonstrated improved performance despite the unbalanced labels. The incorporation of multiple class-related words consistently enhanced performance across all datasets in the unbalanced setting, affirming their utility in ICL.
@@ H.5 Case Study: Inference Cost Evaluation
H.6 Case Study: Sample-Label Visualization
To assess the effectiveness of sample selection based on the discriminative separability of class-related words and the insertion of class-related words into test samples, Fig.9 illustrates the distribution of class-related word logit for selected and other samples, while Fig.10 displays the distribution of inserted class-related word logit under 1-shot ICL learning across testing samples, along with other lass-related word logit distributions under the same conditions.
(a) Logit distribution for selected positive sample
(b) Logit distribution for selected negative sample
(c) Logit distribution for a positive sample
(d) Logit distribution for a negative sample
Figure 9: Visualization of samples’ discriminative separability across class-related words in LLaMA3-8b, evaluated on the SST2 training set using a refined verbalizer. The class-related words are categorized into negative (0-205) and positive (206-429) groups.
(a) Logit distribution for selected positive word
(b) Logit distribution for selected negative word
(c) Logit distribution for a positive word
(d) Logit distribution for a negative word
Figure 10: Visualization of class-related words’ discriminative separability in LLaMA3-8b, evaluated on the SST2 testing set using a refined verbalizer. The samples are categorized into negative (0-911) and positive (912-1820) groups.

### v5

@@ 4.4.2 Effectiveness of Label Organization
In Table 1, inserting multiple class-related words significantly improves ICL performance across all models. To further evaluate the effectiveness of our label organization method, we compare it with a random selection and ordering of multiple class-related words. In Table 5, the lower accuracy observed in the ‘+MLabels-Random’ highlights the advantage of our systematic approach. These results emphasize that thoughtful organization of class-related words is crucial for boosting ICL performance.
@@
Table 5: Accuracy (%) comparison of our multiple class-related words organization method with random multiple words organization in $P_{r}$ (same N as in Table 1).
@@ 4.5 Scalability to Larger LLMs
Table 6 presents the accuracy results of LICL implemented on LLaMA3-8b and LLaMA2-13b, validating the effectiveness and applicability of our method on larger language models. The insertion of multiple class-related words improved accuracy, with LICL achieving a 3.42% increase on LLaMA3-8b and 2.32% on LLaMA2-13b. Notably, for the IMDB dataset on LLaMA2-13b, despite significant sample information loss compared to other methods, our approach maintained strong performance. Details of the evaluation of initial class-related word updates can be found in Appx.G.2.
@@
4.6 General Applicability of Inserted Words
To address potential concerns about overfitting when selecting multiple class-related words from a single dataset, we implement cross-dataset evaluations. We evaluate the efficacy of using two class-related words (N in most binary tasks in our study) selected from the SST2 dataset across the CR and IMDB datasets, as detailed in Table 7.
@@ 4.6 General Applicability of Inserted Words
+MLabels (from same dataset) 95.15 95.60 69.95 73.40
@@
In LLaMA2-7b, performance improvements on the CR may be attributed to similar sub-token/token found in both SST2 and CR (e.g., " unhealthy", " unfair", " good"). No performance decrease is observed in IMDB. In GPT2-xl, despite initial differences in the positive sentiment words across datasets, the class-related words from SST2 still enhanced performance in both CR and IMDB. These results indicate that the selected class-related words exhibit a level of generality, enhancing ICL performance across various datasets with similar labels.
@@ 4.7 Further Analyses
More in-depth analyses are in Appx.H, including pool filtering effectiveness and its ablation study (H.1), label bias evaluation (H.2), the impact of class-related words on predictions whether or not they appear in demonstrations (H.3), the effects of label-balanced demonstration in LICL (H.4), sample-label logit separability visualizations (H.5), and inference cost (H.6).
@@ 5 Conclusion
In this paper, we introduce logit separability, a criterion for measuring the instruction clarity of samples and labels based on their logit value distributions, assessing whether a sample is easily predictable by its true label and whether a label performs well across multiple samples of the same class. To improve ICL with clearer class-related guidance, we propose LICL, a logit separability-based method for selecting and ordering demonstrated samples and labels. Besides, to enrich ICL with broader label information, we present a novel label demonstration method by forming sample-label pairs through the insertion of multiple class-related words, guided by sequential forward search on the logits from selected samples and validation performance. Experimental results confirm that this combined approach significantly enhances ICL classification performance, yielding superior outcomes.
@@ A.2 Label Effectiveness in ICL
A.2.1 Single Class-Related Word as Demonstration Label
This study establishes four distinct label sets for each dataset, utilizing identical samples to form the sample-label pairs in 1-shot ICL, as detailed in Table 8. The reported results represent the average accuracy obtained from five repeated experiments, conducted with seeds 42, 43, 44, 45, and 46 in sample selection during the 1-shot demonstrations. As shown in Fig.5, in LLaMA2-7b, the maximum 49.20% accuracy difference (TREC), and the highest 11.08% (SST2) standard deviation are observed. We also evaluate the label effectiveness in GPT2-xl under the same experimental conditions, with results shown in Fig.6. The results are similar to those in LLaMA2-7b, with a maximum accuracy difference of 48.20% (TREC) and a maximum standard deviation of 10.52% (SST2). These findings suggest that certain class-related words with stronger logit separability, as exampled in Figs.1(b) and 4, can serve as more effective demonstration labels, contributing to greater accuracy and robustness in ICL than predefined class names.
@@ A.2.1 Single Class-Related Word as Demonstration Label
A.2.2 Multiple Class-Related Words as Demonstration Label
@@ Appendix G Enhancement via Initial Class-Related Word Updates
Table 13 presents the effectiveness of initial class-related word updates in 5-shot experiments on the AMAN and TREC datasets, using the same word update information as in the 1-shot setting, as described in Sec.4.2.
@@ H.4 Case Study: Effectiveness of Label Balance in LICL
The impact of label-unbalanced demonstrations varies between the two models. For LLaMA2-7b, a large-size language model, the negative effects of missing label information are less marked, with an average accuracy decrease of 3.37%. The most significant accuracy drops in TREC by 10.40%. The leverage of multiple class-related words mitigated the performance loss, reducing the average accuracy decline to 1.93% in ‘+MLabels_LW’ compared to a label-balanced setting. In contrast, GPT2-xl, a smaller model with 1.5 billion parameters, experienced a marked performance decline under unbalanced conditions, averaging a 12.89% decrease in accuracy. Specifically, SST2 and AGNews experienced over 35% declines. Interestingly, CR demonstrated improved performance despite the unbalanced labels. The incorporation of multiple class-related words consistently enhanced performance across all datasets in the unbalanced setting, affirming their utility in ICL.
H.5 Case Study: Sample-Label Logit Separability Visualization
To assess the effectiveness of sample selection based on the discriminative logit separability of class-related words and the insertion of class-related words into test samples, Fig.9 illustrates the distribution of class-related word logit for selected and other samples, while Fig.10 displays the logit separability of inserted class-related word logit under 1-shot ICL learning across testing samples, along with other lass-related word logit separability under the same conditions.
(a) Logit separability for selected positive sample
(b) Logit separability for selected negative sample
(c) Logit separability for a positive sample
(d) Logit separability for a negative sample
Figure 9: Visualization of samples’ logit separability across class-related words in LLaMA3-8b, evaluated on the SST2 training set using a refined verbalizer. The class-related words are categorized into negative (0-205) and positive (206-429) groups.
(a) Logit separability for selected positive word
(b) Logit separability for selected negative word
(c) Logit separability for a positive word
(d) Logit separability for a negative word
Figure 10: Visualization of class-related words’ logit separability in LLaMA3-8b, evaluated on the SST2 testing set using a refined verbalizer. The samples are categorized into negative (0-911) and positive (912-1820) groups.
H.6 Case Study: Inference Cost Evaluation